OpenStudy (anonymous):

what ratio is not equivalent to 24/36

OpenStudy (destinymasha):

Options?

OpenStudy (anonymous):

A. 3/13 B. 6/24 C. 24/104 D. 36/162

OpenStudy (anonymous):

24/36 = 2/3 3/13 not = 2/3 6/24 = 1/4 not = 2/3 24/104 = 3/13 not = 2/3 36/162 = 2/9 not = 2/3 None of the options are equivalent to 24/36.